McCullough Hospital uses a job-order costing system to assign costs to its patients. Its direct materials include a variety of items such as pharmaceutical drugs, heart valves, artificial hips, and pacemakers. Its direct labor costs (e.g., surgeons, anesthesiologists, radiologists, and nurses) associated with specific surgical procedures and tests are traced to individual patients. All other costs, such as depreciation of medical equipment, insurance, utilities, incidental medical supplies, and the labor costs associated with around-the-clock monitoring of patients are treated as overhead costs.

Historically, McCullough has used one predetermined overhead rate based on the number of patient-days (each night that a patient spends in the hospital counts as one patient-day) to allocate overhead costs to patients. Recently a member of the hospital’s accounting staff has suggested using two predetermined overhead rates (allocated based on the number of patient-days) to improve the accuracy of the costs allocated to patients. The first overhead rate would include all overhead costs within the Intensive Care Unit (ICU) and the second overhead rate would include all Other overhead costs. Information pertaining to the hospital’s estimated number of patient-days, its estimated overhead costs, and two of its patients—Patient A and Patient B—is provided below:

ICU Other Total
Estimated number of patient-days 2,100 10,500 12,600
Estimated fixed overhead cost $ 3,347,400 $ 8,704,500 $ 12,051,900
Estimated variable overhead cost per patient-day $ 257 $ 122
Patient A Patient B
Direct materials $ 4,600 $ 6,300
Direct labor $ 25,250 $ 36,200
Total number of patient-days (including ICU) 10 15
Number of patient-days spent in ICU 0 3

Required:

1. Assuming McCullough uses only one predetermined overhead rate, calculate:

a. The predetermined overhead rate.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

2. Assuming McCullough calculates two overhead rates as recommended by the staff accountant, calculate:

a. The ICU and Other overhead rates.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.
Answer #1
1a
Estimated fixed overhead cost 12051900
Estimated variable overhead cost 1820700 =(2100*257)+(10500*122)
Estimated total overhead cost 13872600
Divide by Total Estimated number of patient-days 12600
The predetermined overhead rate 1101 per patient day
b
Patient A
Direct materials 4600
Direct labor 25250
Applied overhead 11010 =10*1101
Total cost for Patient A 40860
Patient B
Direct materials 6300
Direct labor 36200
Applied overhead 16515 =15*1101
Total cost for Patient B 59015
2a
Estimated fixed overhead cost 3347400
Estimated variable overhead cost 539700 =2100*257
Estimated total overhead cost 3887100
Divide by Estimated number of patient-days 2100
Predetermined ICU overhead rate 1851 per patient day
Estimated fixed overhead cost 8704500
Estimated variable overhead cost 1281000 =10500*122
Estimated total overhead cost 9985500
Divide by Estimated number of patient-days 10500
Predetermined Other overhead rate 951 per patient day
b
Patient A
Direct materials 4600
Direct labor 25250
Applied overhead 9510 =10*951
Total cost for Patient A 39360
Patient B
Direct materials 6300
Direct labor 36200
Applied overhead 16965 =(3*1851)+(12*951)
Total cost for Patient B 59465